OpenStudy (anonymous):

Which of the following is a zero for the function f(x) = (x + 3)(x –7)(x + 5)? (2 points) x = –7 x = –3 x = 3 x = 5

OpenStudy (anonymous):

I said x = 3 because x multiplied 3 times would give ^3 which is a zero for the function

OpenStudy (anonymous):

@mathstudent55 could you help me with this?

OpenStudy (anonymous):

Plug um in, and check.

OpenStudy (anonymous):

I always forget to do that! thank you

OpenStudy (anonymous):

would i have to set the equation equal to 0 or the answer i choose?

OpenStudy (anonymous):

You want f(x)=0. So you don't need to set it equal to zero, just check if it does. and if \[f(-3)\overset{?}{=} 0\] then x=-3 would be a root (or zero) of the equation.

OpenStudy (anonymous):

because -3+3 = 0 and 0 times any multitude of numbers is zero?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

thanks ! could you help me with some more?

OpenStudy (anonymous):

x –8 a factor of the function f(x) = –2x3 + 17x2 –64? Explain. (2 points) Yes. When the function f(x) = –2x3 + 17x2 –64 is divided by x –8, the remainder is zero. Therefore, x –8 is a factor of f(x) = –2x3 + 17x2 –64. No. When the function f(x) = –2x3 + 17x2 –64 is divided by x –8, the remainder is zero. Therefore, x –8 is not a factor of f(x) = –2x3 + 17x2 –64. Yes. When the function f(x) = –2x3 + 17x2 –64 is divided by x –8, the remainder is not zero. Therefore, x –8 is a factor of f(x) = –2x3 + 17x2 –64. I said the first one

OpenStudy (anonymous):

@robtobey ?

OpenStudy (anonymous):

If (x-8) is a factor, then x=8 is a root... So you can sub in to check again. If something is a factor, then there is no remainder. You can check your thinking with numbers, 2 is a factor of 8, and 8/2=4 + zero reminder.